OpenStudy (anonymous):

if I took a full deck of cards, and pulled 1 card out of the deck, then noted the value of the card as a sample (ignoring suits), then took 500 samples and measured the frequency of each card.. How unusual or normal would it be to see a distribution of A = 7.96% K = 6.09% Q = 9.13% 4 = 8.20% 2 = 9.60% How would I go about finding a z score for each result, and an expected standard deviation and variance for the data?

OpenStudy (anonymous):

is 500 samples enough? how could I tell?

OpenStudy (anonymous):

would you need more data? the full data set for the samples is.. A: 7.96% K: 6.09% Q 9.13% J 7.96% T 6.79% 9 9.84% 8 6.32% 7 7.49% 6 7.49% 5 5.62% 4 8.20% 3 7.49% 2 9.60%

OpenStudy (unklerhaukus):

\[\langle x\rangle = \sum xP(x)\]\[\langle x^2\rangle = \sum x^2P(x)\]\[\sigma^2=\langle x^2\rangle-\langle x\rangle^2\]

OpenStudy (unklerhaukus):

you have a table of P(x) values you will need to assign numbers for the x values of J,Q,K, & A, say: 11,12,13, & 14

OpenStudy (unklerhaukus):

\[\langle x\rangle = 14\times6.09\%+13\times7.96\%+\dots \] \[\langle x^2\rangle = 14^2\times6.09\%+13^2\times7.96\%+\dots \]

OpenStudy (unklerhaukus):

What do you get for ⟨x⟩? the average value of the cards? What do you get for ⟨x^2⟩? the average of the square of the value of the cards?

OpenStudy (anonymous):

Thank you Uncle Rhaukaus for your help.. my responses might be a little slow.. Here is what I have so far ... P(x) x P(x) x^2 P(x) 14 7.96% 1.114 15.6016 13 6.09% 0.7917 10.2921 12 9.13% 1.0956 13.1472 11 7.96% 0.8756 9.68316 10 6.79% 0.6790 6.7900 9 9.84% 0.8856 7.9704 8 6.32% 0.5056 4.0448 7 7.49% 0.5243 3.6701 6 7.49% 0.4494 2.6964 5 5.62% 0.2810 1.4050 4 8.20% 0.3280 1.3120 3 7.49% 0.2257 0.6741 2 9.60% 0.1920 0.3840 Sample Mean = 7.9469 x^2 = 77.6193 σ^2 = 77.6193 - 63.1532 = 14.4661 σ = 3.8034

OpenStudy (anonymous):

P(x) x P(x) x^2 P(x) 14 0.07692 1.0769 15.0769 13 0.07692 1.0000 13.0000 12 0.07692 0.9231 11.0769 11 0.07692 0.8462 9.3077 10 0.07692 0.7692 7.6923 9 0.07692 0.6923 6.2308 8 0.07692 0.6154 4.9231 7 0.07692 0.5385 3.7692 6 0.07692 0.4615 2.7692 5 0.07692 0.3846 1.9231 4 0.07692 0.3077 1.2308 3 0.07692 0.2308 0.6923 2 0.07692 0.1538 0.3077 The same logic applied to draws of 1 card out of 13 Sample Mean (x) = 8 sum (x^2) P(x) = 78 σ^2 = 78 - (8^2) = 78 - 64 = 14 σ = 3.7417

OpenStudy (unklerhaukus):

the z-score for each data point is given by z = (x - mean) / σ

OpenStudy (unklerhaukus):

Most of the data should have a z-score less than 1. The first point, A: 7.96%, is near the mean, so it's z score should be very small.